# Implicit Neural Shape Optimization for 3D High-Contrast Electrical Impedance Tomography

### v1

@@ 4 Convergence Analysis
$|d^{2}F(\Gamma;V_{1}(0);V_{2}(0))|\leq C\|V_{1}\|_{C^{3}(\Gamma;\mathbb{R}^{3})}\|V_{2}\|_{C^{3}(\Gamma;\mathbb{R}^{3})}$ (67)
@@
Let $\{z_{k}\}_{k=0}^{\infty}\subset\mathbb{R}^{d}$ be a sequence of latent codes generated by Algorithm 1, where each $z_{k}$ implicitly represents a shape through the neural network $f_{\theta}(z_{k},\cdot)$. Assume that this sequence is uniformly bounded. Let the step size $\alpha_{k}$ be fixed, such that $\alpha_{k}=\bar{\alpha}$ for all $k\in\mathbb{N}$. We also consider the shape derivatives approximated by stochastic gradients $g_{k}$ in the latent space, with the following properties:
@@
    $\nabla\mathcal{L}(z_{k})^{T}\mathbb{E}[g_{k}]\geq\mu\|\nabla\mathcal{L}(z_{k})\|_{2}^{2}\quad\text{and}\quad\|\mathbb{E}[g_{k}]\|_{2}\leq\mu_{G}\|\nabla\mathcal{L}(z_{k})\|_{2}.$ (68)
@@
    $\mathbb{V}[g_{k}]\leq M+M_{V}\|\nabla\mathcal{L}(z_{k})\|_{2}^{2},$ (69)
    where $\mathbb{V}[g_{k}]$ denotes the variance of the stochastic gradient $g_{k}$.
Then, there exists a constant $C$ such that if $0<\bar{\alpha}\leq\min\{C,\frac{\mu}{L\mu_{G}^{2}},\frac{\mu}{LM_{V}}\}$, the following estimate holds:
$\mathbb{E}\left[\frac{1}{K}\sum_{k=1}^{K}\|\nabla\mathcal{L}(z_{k})\|_{2}^{2}\right]\leq\frac{\bar{\alpha}LM}{\mu}+\frac{2(\mathcal{L}(z_{1})-\mathcal{L}_{\inf})}{K\mu\bar{\alpha}}\xrightarrow{K\to\infty}\frac{\bar{\alpha}LM}{\mu},$ (70)
@@
$|d^{2}J(\Gamma;V_{1}(0);V_{2}(0))|\leq C\|V_{1}(0)\|_{C^{3}(\Gamma;\mathbb{R}^{3})}\|V_{2}(0)\|_{C^{3}(\Gamma;\mathbb{R}^{3})}.$ (71)
@@
In our algorithm, the stochastic gradient $g_{k}$ is computed by sampling points on the interface $\Gamma_{z_{k}}$ when approximating the shape derivative integral. This sampling, combined with boundary integral equation (BIE) solutions for the state and adjoint variables, gives us the following properties:
@@
In our neural shape optimization framework, additional constraints arise from the need to preserve topology. Taking these into account, we choose:
$\bar{\alpha}\leq\min\left\{\frac{\mu}{L\mu_{G}^{2}},\frac{\mu}{LM_{V}}\right\}$ (82)
@@
$\mathbb{E}[\mathcal{L}(z_{k+1})|z_{k}]$ $\leq\mathcal{L}(z_{k})-\frac{\bar{\alpha}\mu}{2}\|\nabla\mathcal{L}(z_{k})\|_{2}^{2}+\frac{L\bar{\alpha}^{2}M}{2}$ (83)
@@
Summing over $k=1,2,\ldots,K$:
$\frac{\bar{\alpha}\mu}{2}\sum_{k=1}^{K}\mathbb{E}[\|\nabla\mathcal{L}(z_{k})\|_{2}^{2}]$ $\leq\mathbb{E}[\mathcal{L}(z_{1})]-\mathbb{E}[\mathcal{L}(z_{K+1})]+\frac{KL\bar{\alpha}^{2}M}{2}$ (85)
$\leq\mathcal{L}(z_{1})-\mathcal{L}_{\inf}+\frac{KL\bar{\alpha}^{2}M}{2}$ (86)
@@
$\mathbb{E}\left[\frac{1}{K}\sum_{k=1}^{K}\|\nabla\mathcal{L}(z_{k})\|_{2}^{2}\right]$ $\leq\frac{2(\mathcal{L}(z_{1})-\mathcal{L}_{\inf})}{K\bar{\alpha}\mu}+\frac{\bar{\alpha}LM}{\mu}$ (87)
@@
$\lim_{K\to\infty}\mathbb{E}\left[\frac{1}{K}\sum_{k=1}^{K}\|\nabla\mathcal{L}(z_{k})\|_{2}^{2}\right]$ $\leq\frac{\bar{\alpha}LM}{\mu}$ (88)
@@ 5 Numerical Results
We present comprehensive numerical experiments to validate the effectiveness of our implicit neural shape optimization approach to high-contrast EIT reconstruction in Algorithm 1. The experiments are designed to evaluate both the accuracy of the reconstruction and the robustness of the algorithm under various challenging conditions.
@@ 5.1 Implementation Details
    where $\hat{S}$ and $S^{\dagger}$ are the characteristic functions of the reconstructed and target shapes, respectively.
@@ 5.2.1 Scenario I: Pancreas Reconstruction
Figure 5: Convergence of the algorithm for Scenario I in terms of the loss $\mathcal{L}(z)$ (upper left), the indicator error $e$ (upper left), the volume error (lower left) and the Hausdorff error (lower right).
To evaluate the stability of our method under realistic measurement conditions, we examined the reconstruction performance with significant noise (20%) added to the boundary measurement data. Figure 6 presents the reconstruction results under noisy conditions. Despite the considerable measurement noise, our method successfully recovers the core geometric features of the target shape, preserving both the general curved structure and the approximate proportions. The convergence metrics in Figure 7 further confirm the robustness of our algorithm.
@@
Figure 7: Convergence metrics for the noisy reconstruction (20% noise level): loss function $\mathcal{L}(z)$ (upper left), indicator error (upper right), volume difference (lower left), and Hausdorff distance (lower right).
@@
Figure 9: Convergence of the algorithm for the second example in Scenario I in terms of the loss $\mathcal{L}(z)$ (upper left), the indicator error $e$ (upper right), the volume error $e$ (lower left) and the Hausdorff error (lower right).
@@ 6 Conclusion
This paper presents an implicit neural shape optimization framework for high-contrast EIT reconstruction. Our primary contribution lies in reformulating the shape optimization problem through implicit neural representation, which transforms the original interface reconstruction into an optimization problem in a low-dimensional latent space. This transformation significantly reduces the number of optimization variables while enhancing both computational efficiency and numerical stability. We establish a rigorous theoretical foundation through detailed convergence analysis, providing mathematical guarantees for reconstruction quality under suitable regularity conditions.
Although our approach shows promising results, several aspects need further investigation, including the extension to anisotropic conductivity distributions, adaptation to time-dependent interface evolution, and theoretical analysis of the minimum number of measurements required for stable reconstruction. These directions could further enhance the applicability of the method to more challenging real-world scenarios.
Acknowledgments
This work was partially supported by National Key R&D Program of China 2019YFA0709600, 2019YFA0709602.

### v2

@@ 4 Convergence Analysis
$|d^{2}F(\Gamma;V_{1}(0);V_{2}(0))|\leq C\|V_{1}\|_{C^{2}(\Gamma;\mathbb{R}^{3})}\|V_{2}\|_{C^{2}(\Gamma;\mathbb{R}^{3})}.$ (60)
@@
Let $\{z_{k}\}_{k=0}^{\infty}\subset\mathbb{R}^{d}$ be a sequence of latent codes generated by Algorithm 1, where each $z_{k}$ implicitly represents a shape through the neural network $f_{\theta}(z_{k},\cdot)$. Assume that this sequence is uniformly bounded. Let the step size $\alpha_{k}$ be fixed, such that $\alpha_{k}=\bar{\alpha}$ for all $k\in\mathbb{N}$. We also consider the shape derivatives approximated by stochastic gradients $g_{k}$ in the latent space, with the following properties conditionally on $z_{k}$:
@@
    $\nabla\mathcal{L}(z_{k})^{T}\mathbb{E}[{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}g_{k}|z_{k}}]\geq\mu\|\nabla\mathcal{L}(z_{k})\|_{2}^{2}\quad\text{and}\quad\|\mathbb{E}[g_{k}|z_{k}]\|_{2}\leq\mu_{G}\|\nabla\mathcal{L}(z_{k})\|_{2}.$ (61)
@@
    $\mathbb{V}[{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}g_{k}|z_{k}}]\leq M+M_{V}\|\nabla\mathcal{L}(z_{k})\|_{2}^{2},$ (62)
    where $\mathbb{V}[g_{k}|z_{k}]$ denotes the conditional variance of the stochastic gradient $g_{k}$ given $z_{k}$.
Then, if the step size satisfies $0<\bar{\alpha}\leq\frac{\mu}{L(\mu_{G}^{2}+M_{V})}$, the following estimate holds:
$\mathbb{E}\left[\frac{1}{K}\sum_{k=0}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}K-1}}\|\nabla\mathcal{L}(z_{k})\|_{2}^{2}\right]\leq\frac{\bar{\alpha}LM}{\mu}+\frac{2(\mathcal{L}(z_{0})-\mathcal{L}_{\inf})}{K\mu\bar{\alpha}}\xrightarrow{K\to\infty}\frac{\bar{\alpha}LM}{\mu},$ (63)
@@
Remark.
The assumptions on the stochastic gradient $g_{k}$ are directly motivated by our specific numerical implementation. In Algorithm 1, $g_{k}$ is computed by sampling points on the interface $\Gamma_{z_{k}}$ to approximate the shape derivative integral. This sampling strategy, combined with the boundary integral equation (BIE) solutions for the state and adjoint variables, provides an unbiased estimate of the true gradient given the current latent code $z_{k}$, while inherently introducing a bounded variance.
@@
$|d^{2}J(\Gamma;V_{1}(0);V_{2}(0))|\leq C\|V_{1}(0)\|_{C^{2}(\Gamma;\mathbb{R}^{3})}\|V_{2}(0)\|_{C^{2}(\Gamma;\mathbb{R}^{3})}.$ (64)
@@
Based on the assumptions stated in Theorem 4.4 regarding the conditional expectation and variance of the stochastic gradient $g_{k}$, we obtain the following properties:
@@
Using the step size condition $\bar{\alpha}\leq\frac{\mu}{L(\mu_{G}^{2}+M_{V})}$ assumed in Theorem 4.4, we have
$\mu-\frac{L\bar{\alpha}(\mu_{G}^{2}+M_{V})}{2}\geq\frac{\mu}{2}.$ (75)
@@
$\mathbb{E}[\mathcal{L}(z_{k+1})|z_{k}]\leq\mathcal{L}(z_{k})-\frac{\bar{\alpha}\mu}{2}\|\nabla\mathcal{L}(z_{k})\|_{2}^{2}+\frac{L\bar{\alpha}^{2}M}{2}.$ (76)
@@
Summing over $k=0,1,\ldots,K-1$:
$\frac{\bar{\alpha}\mu}{2}\sum_{k=0}^{K-1}\mathbb{E}[\|\nabla\mathcal{L}(z_{k})\|_{2}^{2}]$ $\leq\mathbb{E}[\mathcal{L}(z_{0})]-\mathbb{E}[\mathcal{L}(z_{K})]+\frac{KL\bar{\alpha}^{2}M}{2}$ (78)
$\leq\mathcal{L}(z_{0})-\mathcal{L}_{\inf}+\frac{KL\bar{\alpha}^{2}M}{2},$ (79)
where we used the fact that the initial point $z_{0}$ is deterministic, hence $\mathbb{E}[\mathcal{L}(z_{0})]=\mathcal{L}(z_{0})$.
@@
$\mathbb{E}\left[\frac{1}{K}\sum_{k=0}^{K-1}\|\nabla\mathcal{L}(z_{k})\|_{2}^{2}\right]$ $\leq\frac{2(\mathcal{L}(z_{0})-\mathcal{L}_{\inf})}{K\bar{\alpha}\mu}+\frac{\bar{\alpha}LM}{\mu}$ (80)
@@
$\lim_{K\to\infty}\mathbb{E}\left[\frac{1}{K}\sum_{k=0}^{K-1}\|\nabla\mathcal{L}(z_{k})\|_{2}^{2}\right]$ $\leq\frac{\bar{\alpha}LM}{\mu}$ (81)
@@ 5 Numerical Results
We present comprehensive numerical experiments to validate the effectiveness of our implicit neural shape optimization approach to high-contrast EIT reconstruction in Algorithm 1. The experiments are designed to evaluate both the accuracy of the reconstruction and the robustness of the algorithm in various challenging scenarios.
@@ 5.1 Implementation Details
    where $\hat{S}$ and $S^{\dagger}$ are the characteristic functions of the reconstructed and target shapes, respectively, and $\{x_{i}\}$ denotes the set of uniformly sampled grid points within the computational domain.
@@ 5.2.1 Scenario I: Pancreas Reconstruction
Figure 5: Convergence of the algorithm for Scenario I in terms of the loss $\mathcal{L}(z)$ (upper left), the indicator error $e$ (upper right), the Hausdorff error (lower left) and the volume error (lower right).
To evaluate the stability of our method under realistic measurement conditions, we examined the reconstruction performance with significant noise (20%) added to the boundary measurement data. Figure 6 presents the reconstruction results with noisy measurements. Despite the considerable measurement noise, our method successfully recovers the core geometric features of the target shape, preserving both the general curved structure and the approximate proportions. The convergence metrics in Figure 7 further confirm the robustness of our algorithm.
@@
Figure 7: Convergence metrics for the noisy reconstruction (20% noise level): loss function $\mathcal{L}(z)$ (upper left), indicator error (upper right), Hausdorff difference (lower left), and volume distance (lower right).
@@
Figure 9: Convergence of the algorithm for the second example in Scenario I in terms of the loss $\mathcal{L}(z)$ (upper left), the indicator error $e$ (upper right), the Hausdorff error $e$ (lower left) and the volume error (lower right).
@@ 6 Conclusion
This work establishes a new paradigm for solving high-dimensional inverse problems by harmonizing the rigor of physics-based modeling with the expressive power of generative AI. We have introduced a “solver-in-the-loop” framework for 3D EIT that fundamentally transforms interface reconstruction from an ill-posed search in infinite-dimensional space to a principled optimization on a compact, learned manifold. By enforcing the governing PDE as a hard constraint through a boundary integral solver while leveraging a differentiable neural prior for geometric regularization, our approach overcomes the fragility of traditional shape optimization and the physical inconsistency often plaguing pure deep learning methods. The theoretical convergence guarantees and robust numerical performance on complex anatomical shapes underscore the reliability of this hybrid architecture.
Looking forward, this framework paves the way for a broader class of scientific machine learning algorithms where data-driven priors and physical laws operate in concert rather than in competition. Immediate extensions include tackling anisotropic conductivities and dynamic interface evolution, while the ultimate vision is to generalize this methodology to other imaging modalities where data is scarce but physical principles are well-understood. Such advancements promise to unlock new capabilities in medical diagnostics, non-destructive testing, and beyond, marking a significant step toward trustworthy, physics-aware AI.
Acknowledgment
The work of J. Chen was partially supported by the National Key R&D Program of China under grants 2019YFA0709600 and 2019YFA0709602. G. Lin would like to acknowledge the support by the National Science Foundation (NSF) under grants DMS-2533878, DMS-2053746, DMS-2134209, ECCS-2328241, CBET-2347401 and OAC-2311848, and by the U.S. Department of Energy (DOE) Office of Science Advanced Scientific Computing Research program under award number DE-SC0023161, and the DOE–Fusion Energy Science program, under grant number: DE-SC0024583.